# TESSERACT: Eliminating Experimental Bias in Malware Classification across Space and Time (Extended Version)

### v1

@@ 8. Discussion
Domain-specific in-the-wild malware percentage . Correctly estimating the malware percentage in dataset is a challenging task and we encourage further measurement studies (Lindorfer et al., 2014; Vadrevu et al., 2013) and data sharing to obtain realistic experimental settings in different security domains.
Correct observation labels. We assume labels in datasets we use are correct. Miller et al. (Miller et al., 2016) found that AVs sometimes change their outcome over time: some goodware may eventually be tagged as malware. However, they also found that VirusTotal detections stabilize after one year; since we are using observations up to Dec 2018, we consider labels we used as reliable. In the future, we may integrate approaches for noisy oracles (Du and Ling, 2010), which select only trustful observations.
Timestamps in the dataset. It is important to consider that some timestamps in a public dataset could be incorrect or invalid. In this paper, we rely on the public AndroZoo dataset maintained at the University of Luxembourg, and we rely on the dex_date attribute as the approximation of an observation timestamp, as recommended by the dataset creators (Allix et al., 2016b). We further verified the reliability of the dex_date attribute by re-downloading VirusTotal (Google, 2004) reports for 25K apps111We downloaded only 25K VirusTotal reports (corresponding to about 10% of our dataset) due to restrictions on our VirusTotal API usage quota. and verifying that the first_seen attribute always matched the dex_date within our time span. In general, we recommend performing some sanitization of a timestamped dataset before performing any analysis on it: if multiple timestamps are available for each object, consider the most reliable timestamp you have access to (e.g., the timestamp recommended by the dataset creators, or the VirusTotal’s first_seen attribute) and discard objects with “impossible” timestamps (e.g., with dates which are either too old or in the future), which may be caused by incorrect parsing or invalid values of some timestamps. To improve trustworthiness of the timestamps, one could verify whether a given object contains time inconsistencies or features not yet available when the app was released (Li et al., 2018). We encourage the community to promptly notify dataset maintainers of any date inconsistencies. In the Tesseract’s project website (section 9), we will maintain an updated list of timestamped datasets publicly available for the security community.
@@
Adversarial ML. Adversarial ML focuses on perturbing training or testing observations to compel a classifier to make incorrect predictions (Biggio and Roli, 2018). Both relate to concepts of robustness and one can characterize adversarial ML as an artificially induced worst-case concept drift scenario. While the adversarial setting remains an open problem, the experimental bias we describe in this work—endemic in Android malware classification—must be addressed prior to realistic evaluations of adversarial mitigations.
9. Availability
@@ A.1. Algorithm Hyperparameters
In adherence to the protocols in (Grosse et al., 2017), DeepDrebin was re-implemented using PyTorch. Initial inputs were aligned with those utilized by Drebin. We mirrored the neural network architecture that demonstrated optimal results as per (Grosse et al., 2017). Yet, we encountered a lack of comprehensive detail on hyperparameter optimization in the source material. Our refined implementation, however, culminated in a marginally elevated score within a 10-fold cross-validation framework (§ 4.4). We postulate that this increment stems from the original research’s optimization for the Accuracy metric (Bishop, 2006), which can be deceptive in imbalanced datasets (Axelsson, 2000)—a common characteristic in the Android domain, dominated by goodware.
A.2. Symbol Table
@@
gw |
Short version of goodware. |
mw |
Short version of malware. |
ML |
Short version of Machine Learning. |
Labeled dataset with malware (mw) and goodware (gw). |
Training dataset. |
Size of the time window of the training set (e.g., 1 year). |
Testing dataset. |
Size of the time window of the testing set (e.g., 2 years). |
Stand for True Positive, False Positive, True Negative, |
and False Negative, respectively |
Size of the test time-slots for time-aware evaluations (e.g., months). |
Performance target of the tuning algorithm in § 4.3; it can be -Score (), Precision () or Recall (). |
AUT(,) |
Area Under Time, a new metric we define to measure performance over time decay and compare different solutions (§ 4.2). It is always computed with respect to a performance function (e.g., -Score) and is the number of time units considered (e.g., 24 months) |
Size of AUT observation time window. |
Estimated percentage of malware (mw) in the wild. |
Percentage of malware (mw) in the training set. |
Percentage of malware (mw) in the testing set. |
Percentage of malware (mw) in the training set, to improve performance on the malware (mw) class (§ 4.3). |
Error rate (§ 4.3). |
Maximum error rate when searching (§ 4.3). |
Model learned after training a classifier. |
Labeling cost. |
Quarantine cost. |
P |
Actual performance; depending on the context, it can refer to AUT with or or . |
A.3. Cumulative Plots for Time Decay
@@
A.4. Tesseract Implementation
Our Python library, Tesseract, incorporates our constraints, metrics, and algorithms, and is designed for seamless integration with standard workflows. Its API, closely aligned with and fully compatible with Scikit-learn (Pedregosa et al., 2011) and PyTorch (Paszke et al., 2019), ensures familiarity for users of these libraries. A detailed overview of Tesseract’s core modules is available in (Pendlebury et al., 2018).

### v2

@@ 8 Discussion
Domain-specific in-the-wild malware percentage . Correctly estimating the malware percentage in dataset is a challenging task and we encourage further measurement studies [42, 69] and data sharing to obtain realistic experimental settings in different security domains.
Correct observation labels. We assume labels in datasets we use are correct. Miller et al. [48] found that AVs sometimes change their outcome over time: some goodware may eventually be tagged as malware. However, they also found that VirusTotal detections stabilize after one year; since we are using observations up to Dec 2018, we consider labels we used as reliable. In the future, we may integrate approaches for noisy oracles [24], which select only trustful observations.
Timestamps in the dataset. It is important to consider that some timestamps in a public dataset could be incorrect or invalid. In this paper, we rely on the public AndroZoo dataset maintained at the University of Luxembourg, and we rely on the dex_date attribute as the approximation of an observation timestamp, as recommended by the dataset creators [3]. We further verified the reliability of the dex_date attribute by re-downloading VirusTotal [31] reports for 25K apps111We downloaded only 25K VirusTotal reports (corresponding to about 10% of our dataset) due to restrictions on our VirusTotal API usage quota. and verifying that the first_seen attribute always matched the dex_date within our time span. In general, we recommend performing some sanitization of a timestamped dataset before performing any analysis on it: if multiple timestamps are available for each object, consider the most reliable timestamp you have access to (e.g., the timestamp recommended by the dataset creators, or the VirusTotal’s first_seen attribute) and discard objects with “impossible” timestamps (e.g., with dates which are either too old or in the future), which may be caused by incorrect parsing or invalid values of some timestamps. To improve trustworthiness of the timestamps, one could verify whether a given object contains time inconsistencies or features not yet available when the app was released [41]. We encourage the community to promptly notify dataset maintainers of any date inconsistencies. In the Tesseract’s project website (section 9), we will maintain an updated list of timestamped datasets publicly available for the security community.
@@
Adversarial ML. Adversarial ML focuses on perturbing training or testing observations to compel a classifier to make incorrect predictions [10]. Both relate to concepts of robustness and one can characterize adversarial ML as an artificially induced worst-case concept drift scenario. While the adversarial setting remains an open problem, the experimental bias we describe in this work—endemic in Android malware classification—must be addressed prior to realistic evaluations of adversarial mitigations.
9 Availability
@@ A.1 Algorithm Hyperparameters
In adherence to the protocols in [33], DeepDrebin was re-implemented using PyTorch. Initial inputs were aligned with those utilized by Drebin. We mirrored the neural network architecture that demonstrated optimal results as per [33]. Yet, we encountered a lack of comprehensive detail on hyperparameter optimization in the source material. Our refined implementation, however, culminated in a marginally elevated score within a 10-fold cross-validation framework (§ 4.4). We postulate that this increment stems from the original research’s optimization for the Accuracy metric [11], which can be deceptive in imbalanced datasets [7]—a common characteristic in the Android domain, dominated by goodware.
A.2 Symbol Table
@@
| gw | Short version of goodware. |
| mw | Short version of malware. |
| ML | Short version of Machine Learning. |
| Labeled dataset with malware (mw) and goodware (gw). | |
| Training dataset. | |
| Size of the time window of the training set (e.g., 1 year). | |
| Testing dataset. | |
| Size of the time window of the testing set (e.g., 2 years). | |
| Stand for True Positive, False Positive, True Negative, | |
| and False Negative, respectively | |
| Size of the test time-slots for time-aware evaluations (e.g., months). | |
| Performance target of the tuning algorithm in § 4.3; it can be -Score (), Precision () or Recall (). | |
| AUT(,) | Area Under Time, a new metric we define to measure performance over time decay and compare different solutions (§ 4.2). It is always computed with respect to a performance function (e.g., -Score) and is the number of time units considered (e.g., 24 months) |
| Size of AUT observation time window. | |
| Estimated percentage of malware (mw) in the wild. | |
| Percentage of malware (mw) in the training set. | |
| Percentage of malware (mw) in the testing set. | |
| Percentage of malware (mw) in the training set, to improve performance on the malware (mw) class (§ 4.3). | |
| Error rate (§ 4.3). | |
| Maximum error rate when searching (§ 4.3). | |
| Model learned after training a classifier. | |
| Labeling cost. | |
| Quarantine cost. | |
| P | Actual performance; depending on the context, it can refer to AUT with or or . |
A.3 Cumulative Plots for Time Decay
@@
A.4 Tesseract Implementation
Our Python library, Tesseract, incorporates our constraints, metrics, and algorithms, and is designed for seamless integration with standard workflows. Its API, closely aligned with and fully compatible with Scikit-learn [52] and PyTorch [51], ensures familiarity for users of these libraries. A detailed overview of Tesseract’s core modules is available in [53].